OpenStudy (anonymous):

1) for the data in the table , does y vary directly with x? If it does, write an equation for the direct variation. X[8, 16, 24] Y[[13, 26, 39] 2) X[32,16,8] Y[24,8,6

OpenStudy (danjs):

If y varies directly with x, then there is some constant or proportionality that relates y to x. y = k*x k = y/x For each y/x, there should be the same constant k.

OpenStudy (danjs):

13/8 = 1.625 26/16 = 1.625 39/24 = 1.625 The first one is a direct proportionality... DO the same for the next part.